The length of a rectangle is 18 m more than 3 times the width. The perimeter is 92 m. Find the length and the width.
Answer by amarjeeth123(569)

Let the width be x.
Then the length is 3x+18.
Perimeter=2(length+width)
2(3x+18+x)=92
4x+18=46
4x=46-18
4x=28
x=7m
Length=39 metres
Width=7 metres